what is the % dissociation of a .15 solution of trimethylamine, pkb=4.20
1
Expert's answer
2020-11-26T05:10:44-0500

α = (kb/c)½

Pkb= 4.20 = -log (kb)

kb= 0.00006309573

α = (0.15/0.00006309573)½

= 48.76